OpenStudy (anonymous):

The three sides of an equilateral triangle measure x + 12, 2x - 5 and 3x - 22 inches. The sides of the triangle measure 17 inches 29 inches 32.5 inches 44.5 inches

OpenStudy (anonymous):

17 inch

OpenStudy (anonymous):

equilateral triangle has 3 same length side. So: x + 12 = 2x - 5 = 3x - 22 x+12=2x-5 x=17 2x-5 = 3x-22 x=17

OpenStudy (anonymous):

thx